3. A monopolist’s demand function is given as Q=2000-10P, where Q the quantity is produced and sold and P is the price per unit in Rwf. If the firm’s marginal cost is 100Rwf
i) Calculate the monopolist’s equilibrium price and quantity’
ii) Suppose the monopolist behaves competitively, how would the answers in (i) above change?
iii) A monopolist is known to cause inefficiency, find the value of the deadweight loss due to the monopoly
1
Expert's answer
2020-12-10T13:41:11-0500

i)Q = 2000 - 10P

At equilibrium

2000 - 10P = 100

10P = 2000 - 100

10P/10 = 1900/10

P = 190

Equilibrium price will be 190Rwf

Equilibrium quantity will be

Q = 2OOO - 10P

Where P = 190

Therefore;

2000 - 10(190) = Q

Q = 2000 - 1900

Q = 100

ii) The price for the commodities will increase or remain the same in the worst scenario.

iii) Deadweight loss = 5* ( P2- P1)*(Q1-Q2)

= 5 x 190 x 100